A student slides a 0.75 kg textbook across a table, and it comes to rest

after traveling 1.2 m. Given that the coefficient of kinetic friction

between the book and the table is 0.34, use the work–kinetic energy

theorem to find the book’s initial speed.


1
Expert's answer
2021-02-18T18:40:04-0500
"\Delta KE= W,""KE_f-KE_i=-F_{fr}d,""0-\dfrac{1}{2}mv_i^2=-\mu mgd,""v_i=\sqrt{2\mu gd}=\sqrt{2\cdot0.34\cdot9.8\ \dfrac{m}{s^2}\cdot1.2\ m}=2.83\ \dfrac{m}{s}."